### v1

# WebAccessVL: Making an Accessible Web via Violation-Conditioned VLM

Abstract
We present a vision-language model (VLM) that automatically edits website HTML to address Web Content Accessibility Guidelines 2 (WCAG2) violations. We formulate this as a supervised image-conditioned program synthesis task, where the model learns to correct HTML given the HTML and its rendering. We collected WebAccessVL, a new dataset with manually corrected accessibility violations, establishing paired training data. We then propose a violation-conditioned VLM that additionally conditions on the WCAG2 violation count to guide the correction process. Experiments demonstrate that our method effectively reduces the average number of violations from 5.34 to 0.44 per website, outperforming commercial LLM APIs (Gemini, GPT-5). A perceptual study confirms that our edited websites maintain the original visual appearance and content.
Figure 1: Given an input HTML with accessibility violations (visualized by boxes in red), e.g., poor contrast, our method makes the HTML WCAG2 compliant by refining it to have improved contrast, a better layout, and an appropriate alt-text.
1 Introduction
An accessible web is “perceivable, operable, understandable, and robust” to everyone, including people with disabilities as defined by WCAG2 [1]. Certain color combinations, such as red and green, may not be distinguished by people with protanopia (color blindness) and should be avoided. Web accessibility ensures that everyone has equal access to information and is legally required in the US (through the ADA [39]) or in the EU (through the EAA [12]). While required by law, surveys by WebAIM [43, 42] found that “95.9% of home pages had detected WCAG2 failures” and that the developers often “lacked the awareness and skills” to design WCAG2 compliant websites. We show that advances in VLMs can help tackle these challenges.
There has been growing interest in leveraging artificial intelligence (AI) to improve web accessibility, such as employing large language models (LLMs) to automatically evaluate accessibility compliance [2]. Other efforts focus on correcting accessibility issues in websites and mobile interfaces by prompting LLMs [17, 30]. These approaches rely exclusively on text-based reasoning without incorporating any visual information. That is, they overlook the rendered appearance of a website, the primary way in which HTML code is perceived by users. Some accessibility violations are more apparent in the rendered view, and less so in the code. This observation motivates us to integrate website renderings as part of the input to the model.
In this work, we formulate the task of making a website accessible as a computer vision problem of image-conditioned program synthesis. Given an HTML document and a screenshot of its rendering, the goal is to produce a modified HTML of the “same website” while ensuring WCAG2 compliance. See Fig. 1 for illustration. The main challenges include (a) designing a model that is aware of the WCAG2 guidelines and applies appropriate modifications; (b) maintaining the output HTML to be faithful to the input website design and rendering; (c) developing a suitable benchmark and metrics to quantify models’ performance.
In this work, we create a dataset consisting of 2,500 webpage HTMLs, of which we manually correct each HTML to better follow WCAG2 standards. Each HTML takes around 7 to 10 minutes to edit. For evaluation, we introduce quantitative metrics to study the accessibility (# of violations) and faithfulness to the original input (similarity in rendering). Finally, we benchmark existing open-API and open-weight LLMs and VLMs.
With paired data, a baseline is to directly perform supervised fine-tuning (SFT), i.e., HTML as input, and generate a more WCAG2-compliant HTML. Differently, we propose to train the model to be violation aware. Unlike SFT, we introduce the number of violations as an additional input to the VLM. As the model is violation-aware, at test-time, we incorporate negative guidance sampling to encourage the model to generate HTML with fewer violations; See Fig. 2 for an overview. Finally, we conduct extensive experiments across 17 models, including LLMs and VLMs, and train our violation-conditioned VLM on open-source models to assess the effectiveness. Our proposed model reduces violations by 73.9% compared with GPT-5, which has the lowest number of violations out of the baselines.
Our main contributions are as follows:
  • We formulate the task of website accessibility as image-conditioned program synthesis and create a dataset (WebAccessVL) of paired HTMLs for training and evaluation on this task.
  • We propose a violation-conditioned VLM, which adapts pre-trained VLMs to improve web accessibility.
  • Through extensive experiments, we show that VLMs outperform LLM for this task and the effectiveness of our approach over baselines.
Figure 2: Overview of our image-conditioned program (HTML) synthesis pipeline. Our model takes in an HTML $\bm{x}$}, the rendered website image $\bm{I}$}, and the violation condition $\bm{c}$}. The Vision-Language Model processes the input tokens and vision tokens extracted from $\bm{I}$} to generate output tokens autoregressively. Our negative guidance sampling refines the token logits to reduce the number of violations in the output, ensuring better compliance with WCAG 2.2 accessibility guidelines.
2 Related Work
AI for accessibility. Existing work has leveraged AI to tackle accessibility challenges, e.g., automated tools designed to suggest source code fixes detected by checkers in mobile apps [30]. Ahmed et al. [2] evaluate ChatGPT’s ability to generate and improve web pages in line with WCAG and show that prompting only resolves simple issues, but struggles with complex tasks. Other works have also employed LLMs, such as ChatGPT, to detect [28] or fix violations through a prompting-based method [33, 3, 17, 5]. In contrast, we introduce a violation-conditioned model that is fine-tuned on a dedicated accessibility dataset created in this work. Importantly, we take both visual and textual input. This allows our method to generate HTML that follows the accessibility guidelines while preserving the original design, providing a more effective solution for web accessibility.
Visual programming and HTML synthesis. Early work by Andreas et al. [4] tackles the problem of visual question answering by assembling neural modules to build an end-to-end differentiable network, one of the first works to output a program from an input image. More recently, VisProg [14] proposes an approach that tackles complex, compositional visual tasks when given natural language instructions and outputs a program that may call other off-the-shelf computer vision models to perform the task, demonstrating impressive zero-shot reasoning capability. Differently, this work focuses on generating accessible HTML while preserving website design.
Research on HTML synthesis primarily focuses on converting rendered website images into code. Beltramelli [7] introduce an end-to-end method for generating code from graphical user interface screenshots using CNNs and RNNs. Building upon VLMs, Lee et al. [22] leverage a pre-trained VLM to transform webpage screenshots into simplified HTML. More recently, the synthetic WebSight dataset [21] provides HTML code paired with corresponding screenshots to facilitate research in this domain, and Roberts et al. [35] introduce a benchmark for evaluating VLMs on extracting structural information from images, including HTML code. Recent studies present datasets and benchmarks for HTML code synthesis using VLMs [49, 45], but these works do not focus on accessibility.
Vision-Language Models have recently achieved remarkable progress by jointly learning from visual and textual data. A typical VLM architecture consists of a visual encoder, a vision-language connector, and a large language model (LLM). For instance, LLaVA [25, 26] is an open-source multimodal framework that leverages the language-only GPT-4 to generate instruction-following data in a multimodal setting and integrates a CLIP-based vision encoder with the LLM.
Next, Qwen VL Series [40, 34] supports multiple languages and can handle multiple images concurrently during training, building on the Qwen language models [46]. Many other deep-net architectures have also been proposed for VLMs. Models such as OtterHD [23], mPLUG-Owl [48], and InternLM-XComposer2-4KHD [9] are designed to utilize high-resolution images, while LLaVA-NeXT [27], Mini-Gemini [24], and MM1 [29] divide the input image into multiple crops.
We show that prompting off-the-shelf VLMs yields some gains in accessibility compliance. In contrast, SFT on our proposed dataset, WebAccessVL, combined with the negative guidance, significantly improves the model’s ability to correct accessibility violations.
3 Approach
Our goal is to build a model that refines the HTML of a website to be WCAG2-compliant while maintaining the website’s content and design. To accomplish this goal, we collected a dataset for SFT (Sec. 3.1). We then propose a violation-conditioned VLM that adapts a pre-trained model for this task (Sec. 3.2).
3.1 WebAccessVL Dataset
To our knowledge, no public datasets are available for training models to refine webpage HTML to follow accessibility guidelines. To address this issue, we collect a dataset and will make it publicly available.
Dataset construction methodology. To construct a set of websites’ HTML, we randomly sampled 2,500 websites from a large-scale HTML dataset provided by Laurençon et al. [21]. During sampling, we verify the availability of essential assets (e.g., images and icons). Websites with missing assets that led to rendering errors were excluded. To ensure reproducibility and prevent similar issues, we save all required assets locally.
To annotate the HTML files, an annotator modifies the code and repeatedly renders each webpage to ensure that its visual design remains consistent with the original version. We use an industrial-grade accessibility checker developed by IBM [19] to minimize accessibility violations to the best of their ability. On average, this process takes approximately 7–10 minutes per webpage. Notably, due to the technical complexity of the task, the annotator requires an advanced degree in computer science.
Dataset statistics. We analyzed the distribution of the collected websites, focusing on the types of accessibility violations and their dependence on visual versus textual information. Among the 26 violation categories, we identified 8 that correspond to vision-related tasks requiring visual understanding. Our analysis shows that 35.8% of violations involve vision-related factors, while the remaining 64.2% are purely language-based, indicating that visual information cannot be overlooked. Below, we briefly describe the major violation types; additional details are provided in Appendix Sec. A1.
Table 1: Accessibility violation statistics.
Violation Percentage (%)
Language violations
     aria_content_in_landmark $25.08$
     html_lang_exists $23.66$
     page_title_exists $23.66$
     skip_main_exists $5.94$
     text_block_heading $5.75$
Vision violations
     text_contrast_sufficient $76.63$
     img_alt_valid $15.90$
     style_color_misuse $6.49$
     caption_track_exists $0.51$
     svg_graphics_labelled $0.47$
Vision violations. In Tab. 1, we present the five most frequent vision-related accessibility issues in the dataset, ranked by their occurrence rates: ① Text contrast sufficient violation means that the ratio of text contrast and background is less than a 3:1 contrast ratio. ② Img alt valid violation is triggered when alternative text is missing. ③ The checker flags a style color misuse violation when there are only color differences to mark required fields, i.e., users who cannot perceive the color will not be able to perform further actions. ④ caption track exists violation occurs when there is no textual information for video content. ⑤ SVG graphics labeled violation is similar to a missing alt text, but it differs in terms of format.
Language violations. As shown in Tab. 1, the top five occurring language-related violations in the dataset are: ❶ aria content in landmark violation is when there is no role or subsection specified. ❷ html lang exists violation is triggered when no language is specified on the website. ❸ page title exists violation is marked if there is no head element and title element in HTML code. ❹ skip main exists violation happens when main tag is missing. ❺ text block heading violation is when there are no headings, i.e., the website has dense blocks of text.
3.2 Violation-conditioned VLM
Problem formulation. Given the input HTML of a website $\bm{x}$} and its rendering $\bm{I}$}, the model $p_{\theta}$ aims to output a revised HTML $\bm{y}$} of the same website but with fewer accessibility violations under WCAG2. To make the model aware of the violations, we introduce a binary indicator ${\bm{c}}\in\{c_{=0},c_{>0}\}$ that specifies whether the refined HTML $\bm{y}$} still violates any WCAG2 guidelines (i.e., whether the number of violations is greater than zero). Conceptually, the model is an image-conditioned program generator. The model is trained by minimizing the negative log-likelihood over the training set $\mathcal{D}$}, i.e.,
$\min_{\theta}\sum_{({\bm{x}},{\bm{I}},{\bm{y}},{\bm{c}})\in{\mathcal{D}}}-\log p_{\theta}({\bm{y}}|{\bm{x}},{\bm{I}},{\bm{c}}).$ (1)
Next, we describe how to adapt $p_{\theta}$ from a pre-trained VLM. An overview of the approach is shown in Fig. 2.
VLM details. Our violation-conditioned VLM is adapted from a pre-trained autoregressive token-based model, where textual and visual inputs are tokenized through text/vision encoders. The model then autoregressively generates a sequence of output tokens, which are subsequently decoded into text.
To prepare the inputs for the VLM, we prepend the input HTML $\bm{x}$} with the prompt “Modify the HTML code to comply with WCAG 2.2 Accessibility Guidelines” and encode it into text tokens ${\bm{t}}=({\bm{t}}_{1},{\bm{t}}_{2},\dots,{\bm{t}}_{L})$. Without the prepended prompt, we found that the models’ outputs are likely not in HTML format. Next, the rendered image $\bm{I}$} is processed by the vision encoder to obtain vision tokens ${\bm{v}}=({\bm{v}}_{1},\dots,{\bm{v}}_{K})$. The violation indicator $\bm{c}$} is tokenized using the text encoder, where $c_{=0}$ is represented with text prompt “The expected output HTML has zero violations” and $c_{>0}$ is represented with “The expected output HTML has non-zero violations”. The target HTML $\bm{y}$} is also encoded into text tokens ${\bm{u}}=({\bm{u}}_{1},{\bm{u}}_{2},\dots,{\bm{u}}_{T})$. SFT of the VLM $\tilde{p}_{\theta}$ is performed by minimizing the token-level negative log-likelihood:
$\min_{\theta}\sum_{({\bm{t}},{\bm{v}},{\bm{u}},{\bm{c}})\in\tilde{{\mathcal{D}}}}\sum_{t=1}^{T}-\log\tilde{p}_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{t}},{\bm{v}},{\bm{c}}),$ (2)
where $\tilde{{\mathcal{D}}}$ denotes the tokenized training set.
Negative guidance sampling. To generate an HTML without violations $\bm{y}$}, we could set ${\bm{c}}=c_{=0}$, sample from $\tilde{p}_{\theta}$, and decode the set of generated tokens. In practice, we found that using a more advanced sampling strategy further reduces the number of violations. Specifically, we incorporate LLM classifier-free guidance with a negative prompt [36, 10, 15, 37]. Let
$\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{c}})\triangleq\log\tilde{p}_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{t}},{\bm{v}},{\bm{c}})$ (3)
denote the conditional logit for the token ${\bm{u}}_{t}$. Negative guidance sampling subtracts a penalty term from the conditional logit, discouraging tokens associated with non-zero violation states.
Formally, the negative guided logit for the token ${\bm{u}}_{t}$ is
$\hskip-28.45274pt\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\triangleq\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{>0})$
$+\gamma\cdot\Bigl(\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{=0})-\ell_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},c_{>0})\Bigr),$ (4)
where $\gamma\in{\mathbb{R}}_{+}$ is the guidance scale. Intuitively, when $\gamma=1$, the sampling reduces to sampling from the conditional distribution with ${\bm{c}}=c_{=0}$. For $\gamma>1$, the logit of tokens associated with a violation greater than zero, i.e., $c_{>0}$, is being downweighted.
Table 2: We show the evaluation metrics using open models and our violation-conditioned VLMs on the test set of WebAccessVL. Within the Baselines, the first block includes open APIs, the second contains open-source LLMs, and the third consists of open-source VLMs.
Base Model VLM # Violation $\downarrow$ Caption-Img CLIP $\uparrow$ Struct. Acc. $\uparrow$ Tree Edit Dist. $\downarrow$
Open API / Weights Baselines Raw Data - 5.335 - - -
GPT-4o 2.929 0.227 0.389 5.081
GPT-5 1.681 0.144 0.005 15.573
Claude3.5 Sonnet 2.310 0.223 0.183 7.603
Gemini 1.5 Flash 6.213 0.229 0.286 5.762
Llama 3.2 3B 4.382 0.173 0.020 14.097
Qwen 2 7B 5.173 0.178 0.022 1.235
Qwen 2.5 7B 5.303 0.134 0.019 0.314
Qwen 2.5 Coder 7B 4.466 0.167 0.022 4.119
GPT-4o Vision 2.918 0.250 0.408 4.939
DeepSeek VL 2 small 5.505 0.149 0.022 2.090
Janus Pro 7B 5.335 0.125 0.022 12.616
Llama 3.2 Vision 11B 4.765 0.165 0.019 15.474
LLaVA 1.5 4.944 0.156 0.021 4.443
LLaVA 1.6 5.203 0.169 0.022 1.095
Qwen 2.5 VL 7B 3.770 0.217 0.019 3.983
Qwen 2 VL 7B 5.214 0.153 0.022 0.125
[…]
The final probability of token ${\bm{u}}_{t}$ is then computed via a Softmax over the vocabulary set $\mathcal{V}$}, defined as:
$\hskip-123.76965pt\tilde{p}^{\text{NG}}_{\theta}({\bm{u}}_{t}\mid{\bm{u}}_{<t},{\bm{t}},{\bm{v}},{\bm{c}})=$
${\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}_{t},\gamma)\Bigr)}\Big/{\sum_{{\bm{u}}^{\prime}\in{\mathcal{V}}}\exp\Bigl(\ell^{\text{NG}}_{\theta}({\bm{u}}^{\prime},\gamma)\Bigr)}.$ (5)
Finally, the sampled output tokens $\bm{u}$} are fed to the text decoder to generate the revised HTML $\bm{y}$}. Note, extra text is occasionally generated outside the HTML structure. Hence, we apply postprocessing to extract the HTML segment using a non-greedy regular expression that matches the content enclosed between the <html> and </html> tags.
4 Experiments
We benchmark various open API and open-weight models, followed by an evaluation of supervised fine-tuning using our proposed dataset. We then conducted a perceptual study to ensure the quantitative metrics aligned with human judgment. Finally, we defer the ablation study to the appendix Sec. A3.
Evaluation. A model is evaluated by its ability to improve accessibility while preserving the original design. We report four metrics to assess the accessibility and visual–structural quality:
  • Accessibility guidance violations $(\downarrow)$: Following IBM’s accessibility checker [19], we report the average number of detected WCAG2 violations per website.
  • Caption-Image CLIP score $(\uparrow)$: We compute the CLIP score between the generated Alt text and the corresponding image to assess the semantic alignment of text and visual content.
  • Structural accuracy $(\uparrow)$: We render both the output HTML and the ground-truth, then compute the Structural Similarity Index Measure [41] (SSIM) to determine if any significant design modifications have occurred. We say a website is structurally consistent if the SSIM value exceeds 0.9.
  • Tree edit distance $(\downarrow)$: We compare the tree edit distance between the raw input HTML and the output HTML to validate HTML design consistency. Leveraging the Document Object Model (DOM) based HTML structure, we consider each HTML tag, such as <div>, <body>, <nav> as a tree node. We set the value of the cost of insert, update, and remove as one to penalize a structural disparity.
Baselines. To evaluate the accessibility quality of model-fixed websites, we selected 16 publicly open models, including both LLMs and VLMs:
  • LLMs: GPT-4o [31], GPT5 [32], Claude3.5 Sonnet [6], Gemini 1.5 Flash [38], Llama3.2 [11], Qwen2 [46], Qwen2.5 [47], and Qwen2.5 Coder [18].
  • VLMs: GPT4o Vision, DeepSeek VL2 [44], Janus Pro [8], Llama3.2 Vision [11], LLaVA1.5 [26], LLaVA1.6 [27], Qwen2 VL [40], and Qwen 2.5 VL [34].
Implementation details. Both LLM and VLM are adapted using LoRA [16]. Across all the models, we use Adam [20] to fine-tune the model for 100 epochs with a learning rate of $2\times 10^{-4}$. During sampling, we do top-1 sampling [13] and set the maximum sequence length to 2048. For negative guidance sampling, the guidance scale is set to 1.1 for LLaVA 1.6 and 1.05 for Llama 3.2 Vision. Input images are resized to $640\times 360$. All experiments fit in a single 48GB Nvidia A6000 GPU. Additional implementation details are provided in Appendix A4.
Figure 3: The percentage improvement in violation for open API/weights models compared to the raw HTML of the test set.
4.1 Comparison with Open API / weights models
To understand the difficulty of this task, we evaluate the performance of publicly available API-based and open-weight models using prompting. For fairness, all models are tested with the same instruction prompt as used in our method (see Appendix A4 for details). In Tab. 2, we report the results across LLMs and VLMs. We also present our fine-tuned models with negative guidance, initialized from LLaVA 1.6 [27] and Llama 3.2 Vision [11].
Results on LLMs. Claude3.5 Sonnet and GPT-4o are the best performing open APIs. They achieve moderate gains in improving accessibility with some degree of content/design preservation. While GPT-5 achieves the lowest number of violations, it significantly alters the website design, as indicated by its low structure accuracy and high tree edit distance. For Gemini 1.5 Flash [38], it slightly increased the number of violations. When comparing to open-source LLMs, such as Llama 3.2 or Qwen, we find that they, in general, perform worse than the open-API models.
Results on VLMs. For models that take in visual inputs, we further provide the rendering of the HTML. For GPT-4o Vision, it achieved a lower number of violations compared to its pure LLM counterpart and achieved an improved structure accuracy. Other models are able to fix some violations, except DeepSeek VL2 small [44] that slightly increases the violation count. We also visualized the percentage improvement in violation from the raw data in Fig. 3.
Results on Ours. As shown in Tab. 2, our method convincingly outperforms the baseline models. Our approach significantly lowers the number of violations to 0.5 violations per website on average. Our method also produces more accurate alt text and better preserves the website structure. Our method on Llama 3.2 Vision and LLaVA 1.6 shows an improvement in the number of violations of 91.8% and 90.5%. For structural accuracy, our models outperform all open-source models. The slightly lower SSIM compared to GPT-4o is due to adjustments of text contrast, which may reduce the SSIM and tree edit distance but improve accessibility significantly. Overall, these results demonstrate the importance of the proposed WebAccessVL dataset with SFT for developing automated systems to improve accessibility. We now study SFT in more detail.
Figure 4: The percentage improvement in violation for SFT models compared to the raw HTML of the test set.
Figure 5: The percentage of vision/language violations out of all the violations from each LLM and VLM.
Table 3: Results of SFT the violation-conditioned VLM with different pre-trained models on WebAccessVL. The results of LLMs are in the first block, and those of VLMs are in the second block.
[…]
SFT Models Llama 3.2 1.119 0.208 0.255 3.104
Qwen 2 7B 0.896 0.217 0.355 2.776
Qwen 2.5 7B 0.743 0.220 0.310 2.926
Qwen 2.5 Coder 7B 0.579 0.221 0.304 2.947
LLaVA 1.5 0.810 0.244 0.313 2.989
LLaVA 1.6 0.527 0.262 0.307 3.049
Qwen2 VL 0.598 0.220 0.317 2.938
Qwen2.5 VL 0.700 0.261 0.330 2.825
Llama 3.2 Vision 0.451 0.260 0.331 3.004
[…]
4.2 Comparison with SFT models
We examine the effect of supervised fine-tuning (SFT) on our violation-conditioned model with different pre-trained models. Results are shown in Tab. 3, where we observe that all models fine-tuned on WebAccessVL show a reduction in the number of violations.
Importance of visual information in SFT. In Fig. 4, we plot the percentage of improvement of violation # relative to the raw data. We observe that on average, VLMs tend to improve more over LLMs. For instance, the Llama 3.2 model with and without vision has a large gap in the percentage of improvement. From Tab. 3, we see that Llama 3.2 (LLM) has an average of 1.119 violations per HTML code, which drops to 0.451 when visual input is added. Furthermore, the caption-image CLIP score increases from 0.208 to 0.260, highlighting the benefit of visual information. In other words, the VLM variant better preserves the rendered structure, as indicated by higher structural accuracy and a lower tree edit distance. A similar trend is also observed between Qwen 2.5 7B and its VLM variant.
Analysis on vision vs. language violations. To further investigate whether violation-conditioned VLMs outperform LLMs, we plot the proportion of vision and language violations relative to the total violations made by each model in Fig. 5. Across the same backbone, we observe that Qwen 2.5 VL exhibits over 25% fewer vision-related violations than Qwen 2.5. This decrease is primarily due to improved handling of text contrast and image captioning in alt attributes. On average, 54.5% of the violations for LLMs are related to vision tasks, compared to 46.3% for VLMs. This reduction of 8.1% demonstrates that VLMs with visual input are more effective at addressing vision violations than pure LLMs. Critically, LLMs often fail to detect insufficient text contrast, as they rely solely on textual HTML content and rarely pick up on visual styles such as foreground-background color pairs. We hypothesize that as VLMs can directly inspect rendered screenshots, this enables them to correct vision violations more reliably.
Ablation of negative guidance. In Tab. 3, we report results with and without negative guidance. We observe that incorporating negative guidance into our violation-conditioned VLMs further reduces the number of violations compared to their counterparts without negative guidance. Specifically, the violation count for LLaVA 1.6 decreases from 0.527 to 0.509 with standard deviation of 0.0012, and for Llama 3.2 Vision, it drops from 0.451 to 0.439 with standard deviation of 0.0017, both averaged over three random seeds. At the same time, the other metrics remain comparable to those of the version without negative guidance.
[…]
Refer to caption Refer to caption Refer $\columncolormygrayRefer to caption$
Refer to caption Refer to caption Refer $\columncolormygrayRefer to caption$
Figure 6: Qualitative results with Ours, Claude 3.5 Sonnet [6], and ChatGPT4o Vision [31]. Yellow arrows highlight the fixed components, and red arrows show inaccessible parts.
4.3 Qualitative Results
In Fig. 6, we present qualitative comparisons on in-the-wild websites to show our model’s ability to generalize beyond the WebAccessVLdataset. In the first row, we observe that our model fixed the header, main background image, button, and footer contrast and put a gray bar at the bottom. However, Claude 3.5 did not change the header contrast but did change the gray bar at the bottom to enhance readability. GPT-4o generated a text close to the button in the main image, which makes it hard to read, and removed one of the products.
In the second example, our model fixed the contrast of the header, top three icons, green buttons, and the footer to improve readability. In contrast, Claude 3.5 did not fix the contrast of the header. Also, the footer color has insufficient contrast. GPT-4o missed two products and failed to put the footer in the right location. Overall, our model can fix contrasts of visual assets to improve visibility and make minimal changes to the website structure/style.
4.4 Perceptual Study
Study design. We conducted a perceptual study to assess how well the modified HTML files preserve both the website’s content and visual style. We compare against Claude 3.5 and GPT-4o, which achieved a relatively low number of accessibility violations while maintaining overall design fidelity, as observed in Tab. 2. Although GPT-5 achieves a low violation count, it makes large changes to the website’s design.
Each participant is shown the reference webpage along with outputs from our method and two baselines, Claude 3.5 and GPT-4. Specifically, participants were asked, “Which website (A, B, or C) best maintains the content and visual style of the reference image? In the event of a tie, please select the one with the better visual design.” Participants may select only one option for each question or leave the question empty. We gathered statistics on 13 questions and received responses from 31 participants.
Statistical hypothesis setup. We analyze user preference data using a chi-squared test of independence to determine whether the distribution of selections differs significantly across the three methods. Let the three methods be denoted as $A$, $B$, and $C$ (AA is Ours, $B$ is Claude 3.5, and $C$ is GPT-4o). We first perform a chi-squared test of independence on the following hypothesis (More details in Sec. A5):
$H_{0}^{\mathrm{global}}:$ $\;p_{A}=p_{B}=p_{C},$ (6)
$H_{1}^{\mathrm{global}}:$ $\;\text{at least one }p_{i}\neq p_{j}.$ (7)
If $H_{0}^{\mathrm{global}}$ is rejected at level $\alpha=0.05$, we proceed to post-hoc pairwise comparisons between each pair $(i,j)\in\{(A,B),(A,C)\}$. For each pair, we test
$H_{0}^{\,i>j}:p_{i}=p_{j},\qquad H_{1}^{\,i>j}:p_{i}>p_{j}.$ (8)
We apply the Bonferroni correction to control the family-wise error rate over the three comparisons. The adjusted significance threshold is $\alpha_{\mathrm{adj}}=\frac{0.05}{3}\approx 0.0167.$
Results and statistical analysis. We aggregate selections over all 13 questions with 401 valid responses, with counts:
$(n_{A},n_{B},n_{C})=(248,\;54,\;99).$ (9)
The chi-squared test of independence on the 3×1 contingency table yielded
$\chi^{2}_{\mathrm{global}}(2,\,N=401)=154.27,\quad p<0.001.$ (10)
Since $H_{0}^{\mathrm{global}}$ is rejected at $\alpha=0.05$, we perform post-hoc pairwise chi-squared tests with Bonferroni correction. The one-tailed comparisons for our method $A$ against $B$ and $C$ yield:
$\chi^{2}_{A>B}(1,\,N_{A,B}=302)=130.01,\quad p_{A>B}<10^{-16},$ (11)
$\chi^{2}_{A>C}(1,\,N_{A,C}=347)=64.00,\quad p_{A>C}<10^{-15}.$ (12)
Both $p_{A>B}$ and $p_{A>C}$ are below $\alpha_{\mathrm{adj}}=0.0167$. These pairwise results demonstrate that participants significantly selected our method over both Claude 3.5 and GPT-4o, i.e., our method better preserves the website’s content and visual style. This conclusion is in agreement with the quantitative structural metrics (Struct. Acc and Tree Edit Dist.) presented in Tab. 2.
5 Conclusion
In this work, we formulate the task of improving a website’s accessibility as an image-conditioned program synthesis problem. We construct a dataset consisting of paired HTML files that have been manually modified to comply with WCAG2. We propose training VLMs in a violation-conditioned manner to address this task. Empirical results show that our method reduces the average number of violations from 5.34 to 0.44 per website on the test set. Compared to GPT-5, our approach achieves a 73.9% reduction in violations. We also show that LLMs are suboptimal in understanding visual content when relying solely on HTML code. This work is a first step toward automating web accessibility. To support further research, we will release both the code and the dataset.
Appendix
The appendix is organized as follows:
  • We describe the details of each accessibility violation in Sec. A1.
  • We provide additional qualitative results in Sec. A2. Also see the attached HTML files in the supplementary materials.
  • We conduct more ablation studies on the training module and training set size in Sec. A3.
  • Implementation, experimentation details, and additional prompt experiments are provided in Sec. A4. We will release the code and dataset.
  • We explain the test statistic for the user study in Sec. A5.
  • We report the training and inference times of our violation-conditioned VLM in Sec. A6.
  • We document how we annotated websites to improve WCAG2 compliance for the WebAccessVL dataset in Sec. A7.
A1 Dataset Violation Details
Vision violations. Below, we describe the top-five vision violations:
 ① Text contrast sufficient violation means that the ratio of text contrast and background is less than a 3 to 1 contrast ratio. This violation requires visual understanding to fix the color contrast while maintaining the website’s design consistency. Naively using a white background and black text would resolve this violation, but it would also lead to large changes in the website design.
 ② Img alt valid violation is triggered when there is a missing alternative text. Each visual representation, such as a product image, logo, or any image, is required to have a short description of the image. For example, the checker verifies that the <img> tag has an alt attribute or not in the HTML. Vision information is needed as the description needs to match the image content.
 ③ The checker flags a style color misuse violation when there are only color differences to mark required fields, i.e., users who cannot perceive the color will not be able to perform further actions. This violation can be fixed using a visual indication such as the asterisk (*) in the required field. Note that the color needs to be chosen carefully; an inappropriate color could recursively trigger color contrast violations.
 ④ caption track exists violation occurs when there is no textual information on a given video content. This violation can be solved by providing captions containing textual information on the video content.
 ⑤ svg graphics labeled violation is similar to a missing alt text, but it differs in terms of format. SVG can have multiple layers of components, and each component should have its corresponding label, i.e., labeling the meaning for all the components. This additional information can be provided using an accessible specific tag to assistive technologies using aria-label; ARIA stands for Accessible Rich Internet Applications.
In summary, these vision violations need some degree of reasoning over the visual content of the website and would be difficult to address using only language information.
Language violations. We now describe the top-five language violations:
 ❶ aria content in landmark violation is when there is no role or subsection specified. Without having the role, screen readers or other assistive technologies cannot understand the elements on the web page. A textual understanding of content is required to assign a proper role from the eight possible roles, using the role tag.
 ❷ html lang exists violation triggers when there is no specified type of language on the website. It is necessary to support assistive technology such as speech synthesizers to adapt the pronunciation to the specific language.
 ❸ page title exists violation is marked if there is no head element and title element in HTML codes. A page title is important because it is the first part that assistive technology accesses the website and tells the users so that they can confirm they are accessing the right website.
 ❹ skip main exists violation happens when main tag is missing. Without the tag, users with assistive technologies cannot directly go to the main elements, and users need to go through all unnecessary content, such as banners or navigation elements.
 ❺ text block heading violation is when there are no headings, i.e., the website has dense blocks of text. With proper headings, they allow users with assistive technology to go through blocks by accessing a summary of contents and quickly navigating to other sections.
Overall, these language violations are mainly related to the HTML structuring and summarizing of the website’s text content. Addressing these violations will not involve the usage of vision information.
A2 Additional Qualitative Results
We show additional qualitative results using our violation-conditioned model, Claude 3.5 Sonnet [6], and ChatGPT4o Vision [31] in Fig. A1 to Fig. A6.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A1: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours modifies the contrast between the header, a button, and the descriptions in each card, while Claude 3.5 fails to put valid image links. Moreover, ChatGPT4o could not correct the contrast of the header. Please zoom in for details.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A2: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours corrects the text contrast in the navigation bar, descriptions in each image, and side bars. Claude 3.5 does not have appropriate contrast on the title of each card and is missing the thumbnails. ChatGPT4o also has not completely fixed the text contrast in the left side bar, and generates a different website layout ratio. Please zoom in for details.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A3: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours successfully put an accessible contrast in buttons and the texts. Claude 3.5 and chatGPT4o failed to put an appropriate color contrast on the buttons. Please zoom in for details.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A4: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows that it generates an accessible contrast on the header, description, and footer. Moreover, ours preserves the website style by putting all four images in a row, while Claude 3.5 and chatGPT4o could not preserve the original design. Please zoom in for details.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A5: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it can keep the style while correcting inaccessible elements by changing their color contrast values. Claude 3.5 failed to preserve the original web design by making thicker header and footer thicker. Also, it could not put the correct image URLs. Please zoom in for details.
Input Claude3.5
Refer to caption Refer to caption
ChatGPT4o Ours
Refer to caption Refer to caption
Figure A6: Yellow arrows highlight the fixed components, and red arrows show inaccessible parts. Ours shows it fixes text contrast in the header, image cards, and the footer by following the original web style. Claude 3.5 did not preserve the original style by making larger selection boxes and missing images. ChatGPT4o failed to fix the proper text and button contrasts to comply with the accessibility guideline. Please zoom in for details.
A3 Ablation Study
[…]
Qwen2 VL QV 1.123 (79.0) 0.202 0.318 2.893
QKVO 1.127 (78.9) 0.243 0.309 3.101
QKVO+ FC + LN 0.621 (88.4) 0.255 0.325 2.794
QKVO + VP + VR 0.598 (88.8) 0.220 0.317 2.938
Qwen2.5 VL QV 3.491 (34.6) 0.164 0.632 1.554
QVKO 1.625 (69.6) 0.220 0.382 3.210
QKVO + FC + LN 1.672 (68.7) 0.229 0.359 2.198
QKVO + VP + VR 0.700 (86.9) 0.261 0.330 2.825
Llama 3.2 Vision QV 0.548 (89.7) 0.249 0.295 2.995
QKVO 0.582 (89.1) 0.242 0.302 2.996
QKVO + MLP 0.885 (83.4) 0.260 0.356 2.884
QV + DP + UP + GateP 0.451 (91.6) 0.260 0.331 3.004
LLaVA 1.5 QV 0.958 (82.0) 0.234 0.326 2.940
QVKO 1.349 (74.7) 0.247 0.395 3.076
QKVO + Vision proj 0.810 (84.8) 0.244 0.313 2.989
QKVO+ FC + LN 1.182 (77.8) 0.232 0.450 2.488
LLaVA 1.6 QV 0.627 (88.2) 0.248 0.310 2.989
QVKO 0.510 (90.4) 0.259 0.293 3.097
QKVO + Vision proj 0.618 (88.4) 0.246 0.311 3.032
QKVO+ FC + LN 0.527 (90.1) 0.262 0.307 3.049
Table A1: Results of ablating different training modules from the test set of WebAccessVL.
Ablating trainable modules for SFT VLMs. We conduct ablation studies to analyze the impact of different training modules on our model’s performance. In Tab. A1, we present results across multiple open-source VLMs and their training module variations. The results indicate that Query, Key, Value, and Out Projection generally enhance Caption-Image CLIP similarity but may increase violations. Incorporating fully connected layers, layer norms, or structured projections helps reduce violations while maintaining alignment. Among the variations, we report the one with the fewest violations in Tab. 3.
Ablating training sample size for SFT VLMs. We investigate the effect of training sample size in Tab. A2. As the number of training samples increases, violations generally decrease, though not strictly, suggesting that models can achieve better alignment with accessibility guidelines when trained with more appropriate data. For Caption-Image CLIP similarity, we observe an increasing trend in Llama 3.2 Vision [11] and LLaVA 1.5 [26], while other models do not exhibit a consistent trend.
[…]
Qwen2 VL 100 1.364 (74.4) 0.217 0.287 2.835
200 0.903 (83.1) 0.221 0.304 2.925
300 0.836 (84.3) 0.209 0.289 3.034
400 0.888 (83.4) 0.208 0.316 2.977
Qwen2.5 VL 100 2.153 (59.6) 0.212 0.414 2.182
200 3.054 (42.8) 0.162 0.612 1.517
300 3.276 (38.6) 0.137 0.666 1.224
400 2.102 (60.6) 0.188 0.479 2.337
Llama 3.2 Vision 100 0.674 (87.4) 0.227 0.237 2.989
200 0.602 (88.7) 0.243 0.320 2.847
300 0.574 (89.2) 0.246 0.328 2.967
400 0.589 (89.0) 0.260 0.318 3.010
LLaVA 1.5 100 1.065 (80.0) 0.234 0.253 3.786
200 1.060 (80.1) 0.239 0.286 3.618
300 1.278 (76.0) 0.231 0.362 3.168
400 0.870 (83.7) 0.242 0.315 3.025
LLaVA 1.6 100 0.895 (83.2) 0.211 0.308 3.135
200 0.950 (82.2) 0.151 0.290 3.101
300 0.698 (86.9) 0.226 0.320 3.047
400 0.597 (88.8) 0.202 0.317 2.977
Table A2: Results of ablating different training sizes from the test set of WebAccessVL.
A4 Implementation Details & Prompt Ablation
Model Trainable Module Trainable %
Qwen 2 VL QKVO, Vision Proj, Vision Resampler 6.08
Qwen 2.5 VL QKVO, Vision Proj, Vision Resampler 6.08
Llama 3.2 Vision QV, Down, Up, Gate Proj 22.06
LLaVA 1.5 QKVO, Vision Proj 13.90
LLaVA 1.6 QKVO, FC1&2, Layer Norm 10.90
Table A3: Trainable module and trainable parameter percentage of Violation-conditioned VLM.
Training details. The trainable module and the trainable parameter percentage of each model are provided in Tab. A3. Across all models, we train LoRA [16] with a rank of $r=8$ and $\alpha=16$, resulting in a scaling factor of $\alpha/r=2$. The dropout rate is set to 0.05. Due to memory constraints, we use gradient accumulation to enable training with a larger batch size. Specifically, we train Llama 3.2 Vision [11] with a batch size of 32 and LLaVA 1.6 [27] with a batch size of 24, both using negative guidance. The number of epochs for all the training runs is 100. During training, the input text prompt is:
[frame=single]
Modify the following HTML code to comply with WCAG 2.2 accessibility
guidelines. Ensure the output is a properly formatted HTML file without
additional explanations or descriptions. Return only the modified HTML
code without extra text or commentary.
The expected output HTML has <zero/non-zero> violations.
Input HTML:
<input HTML x>
Inference details. During inference, we generate the output HTML code by using the same prompt as training. However, in order to conduct negative-guided sampling, we performed two forward passes, using $c_{=0}$ and $c_{>0}$, i.e., putting zero and non-zero in the condition.
API calls. We use a public API from OpenAI for ChatGPT4o with an image to run an inference using the following code as VLM. For LLM, we did not send the image data as the official document introduced. For all publicly available models, including Gemini 1.5 [38] and Claude 3.5 [6], we use the same prompt but different API calls using their API functions.
import base64
from openai import OpenAI
# Function to encode the image
def encode_image(image_path):
with open(image_path, "rb") as image_file:
return base64.b64encode(image_file.read()).decode("utf-8")
client = OpenAI()
# HTML = HTML Path.
# image_path = A rendering from the HTML.
base64_image = encode_image(image_path)
with open(html, "r", encoding="utf-8") as file:
html_content = file.read()
response = client.chat.completions.create(
model="gpt-4o",
messages=[
"role": "user",
"content": [
"type": "text",
"text": f"ModifythefollowingHTMLcodetocomplywith
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣accessibilityguidelines,WCAG2.2,andonlyoutputthe
␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣␣HTMLcodewithoutexplanations.{html_content}",
"type": "image_url",
"image_url": {"url": f"data:image/jpeg;base64,{base64_image}"},
Refined prompt experiments. We further examine whether a higher-quality prompt can improve the performance of baseline models. Following the OpenAI Prompt Engineering Guide, we redesigned the prompt to include explicit context, detailed instructions, and concrete accessibility requirements. The full prompt is:
Context:
You are a professional software engineer specializing in improving web
accessibility and WCAG 2.1 compliance. You are tasked with auditing and
fixing HTML source code to ensure it meets accessibility standards.
Instructions:
- Review and apply your comprehensive knowledge of WCAG 2.1.
- Fix accessibility violations in the provided HTML code.
- Maintain the original visual and interactive design as closely as possible.
- Prioritize interactive elements that require user input or navigation.
- Ensure semantic HTML is used correctly (e.g., headings hierarchy,
  <main>, <nav>, <section>, etc.).
- Only add ARIA roles or attributes when necessary.
- All <img> tags must have descriptive alt attributes.
- Ensure color contrast meets WCAG AA.
- All interactive elements must be keyboard accessible,
  focusable, and correctly labeled.
- Provide <label> elements for form inputs.
- Ensure <html> has a correct lang attribute and include a meaningful <title>.
{HTML Code goes here}
Our previous prompt simply stated:
Modify the HTML code to comply with WCAG 2.2 Accessibility Guidelines.
{HTML Code goes here}
We evaluate both prompts using the same test set and report the number of accessibility violations after correction. As shown in Tab. A4, the refined prompt improves the performance of baseline models. For example, ChatGPT-4o reduces violations from 2.929 to 1.793. However, our method remains substantially more effective, achieving only 0.439 violations.
Model Number of Violations
ChatGPT-4o (New prompt) 1.793
ChatGPT-4o (Old prompt) 2.929
Gemini-2.0-flash (New prompt) 2.134
Gemini-1.5-flash (Old prompt) 6.213
Ours 0.439
Table A4: Remaining accessibility violations after correction under different prompt designs.
Impact of visual input. To verify that the visual input is not ignored by the model, we conduct an experiment in which the rendered image is replaced with a solid black image while leaving the HTML code unchanged. Under this setting, LLaVA 1.6 shows a clear drop in performance, with violations rising from 0.527 to 0.832, a 56.9% increase. This indicates that the model does use visual cues to detect and fix accessibility errors.
A5 Test Statistics of User Study
Let $n$ be the total number of valid votes (skipped answers are excluded), and let $n_{i}$ be the number of votes for method $i\in\{A,B,C\}$. Under $H_{0}^{\mathrm{global}}:p_{A}=p_{B}=p_{C}$, the expected count for each method is
$E_{i}\;=\;n\,p_{0},\qquad p_{0}=\tfrac{1}{3}.$
The Pearson chi-squared statistic is
$\chi^{2}\;=\;\sum_{i\in\{A,B,C\}}\frac{(n_{i}-E_{i})^{2}}{E_{i}}\;=\;\sum_{i\in\{A,B,C\}}\frac{(n_{i}-n/3)^{2}}{\,n/3\,}$,
which follows a $\chi^{2}$ distribution with $2$ degrees of freedom under $H_{0}^{\mathrm{global}}$. We report the associated $p$-value and Cramér’s $V=\sqrt{\chi^{2}/(n\,(k-1))}$ with $k=3$ as an effect size.
A6 Training and Inference Time
We report the training and inference time of the supervised fine-tuned VLMs in Tab. A5. Specifically, we provide the training time for one epoch and the average inference time for a single HTML input. As expected, the training time remains unchanged with or without our violation-conditioning technique. Additionally, the inference time matches the corresponding pre-trained models, as our approach does not introduce new model parameters.
Model Training Time (sec.) Inference Time (sec.)
Qwen 2 VL 168.6 3.08
Qwen 2.5 VL 187.8 3.20
Llama 3.2 Vision 368.4 7.32
+ Neg. Guidance - 9.40
LLaVA 1.5 144.0 10.50
LLaVA 1.6 202.8 5.64
+ Neg. Guidance - 9.30
Table A5: Training time of VLMs fine-tuned on WebAccessVL for one epoch and average inference time for one HTML code.
A7 Website Correction Process for WebAccessVL
We introduce how we manually correct a website to comply with WCAG guidelines, as we introduced in Sec. 3.1. First, we load the HTML code and visually check the content to preserve its design style. We run the IBM checker [19] on the HTML code to check violations using the following command:
      achecker --outputFolder output myHTML.html      It will output violations, if any, and their XPath to locate in the HTML code, along with their snippet. It also includes an IBM website link for a brief solution and its message to explain the violation, like the following:
      Level: violation    XPath: /html[1]/body[1]/div[1]/div[2]/table[1]/tbody[1]/tr[2]/td[3]    Snippet: <td aria-colindex="2" role="cell">    Help: https://able.ibm.com/rules/archives/...    - Message: An explicit ARIA ’role’ is not valid for <tr> element     within an ARIA role ’table’ per the ARIA in HTML specification      Next, we go to the XPath that is flagged as the violation, and we fix it to minimize the changes in the original design. We then visually inspect the updated website to ensure there are no significant structural or color changes. While reducing the number of violations is important, preserving the original design remains a key priority. We repeat this process until all violations are resolved or until the remaining violations require significant modifications to the existing design. On average, fixing a single website takes 7 to 10 minutes. Constructing WebAccessVL, which contains 2500 manually corrected websites, therefore requiring approximately 417 hours in total.
A8 Broader Impact
This work contributes to improving web accessibility by using vision-language models to automatically reduce WCAG2 violations in HTML. It has the potential to benefit users with disabilities and support developers by lowering the technical barrier to creating accessible websites. By relying on open-weight models, our approach also offers a privacy-conscious alternative to commercial services. However, automated fixes may optimize for rule-based compliance without ensuring real usability, and over-reliance on such systems could reduce the incentive for inclusive design practices. We view this method as a tool to support, not replace, human developers and encourage continued collaboration with accessibility experts and users.